The difference between two 2 -digit numbers is 46. What might the numbers be? Find as many answers as you can. Write a subtraction equation for each answer.
Answer by greenestamps(13327)



Are you kidding? What is the point in writing out all the solutions as subtraction equations....?!

56-10 = 46
57-11 = 46
58-12 = 46
...
98-52 = 46
99-53 = 46